OpenStudy (anonymous):

When you see a traffic light turn red, you apply the brakes until you come to a stop. Suppose your initial speed was 11.2 m/s, and you come to rest in 34.7 m. How much time does this take? Assume constant deceleration.

OpenStudy (anonymous):

when we come to rest final velocity is zero using S=ut+0.5*a*t*t =ut+0.5*((vf-vi)/t)*t*t since due to constant deceleration =ut-0.5*t =0.5*ut 34.7=0.5*11.2*t t=34.7/(0.5*11.2) =34.7/5.6 =6.196 seconds